Two wires, both carrying current out of the page have a current of magnitude 2.0 mA and 3.0 mA, respectively. The first wire is located at (0.0 cm 5.0 cm) while the other wire is located at (12.0 cm 0.0 cm). What is the magnitude of the magnetic force per unit length of the first wire on the second and the second wire on the first.


1
Expert's answer
2021-05-26T15:32:45-0400

"F_{12}=F_{21}=\frac{\mu_0}{4\pi}\cdot\frac{2I_1I_2}{d}\cdot l"


"l=1\ (m)"


"d=\sqrt{5^2+12^2}=13\ (cm)=0.13\ (m)"


"F_{12}=F_{21}=\frac{4\pi\cdot10^{-7}}{4\pi}\cdot\frac{2\cdot 2\cdot10^{-3}\cdot3\cdot10^{-3}}{0.13}\cdot 1=3.1\cdot10^{-12}\ (N)=3.1\ (pN)" . Answer